Please help me solve the following problem. I have tried everything I can think of and am completely lost. the problem is: You want to install a 3 yard wide walkway around a circular swimming pool. The diameter of the pool is 18 yards what is the area of the walkway? I tried the area formula and subtracting the area of the pool but that was not it. Can you also show me the formula you used as well?
Thank you very very much,
Kayla

Answer by jim_thompson5910(35256)

Area of Pool+Walkway

A = pi*r^2

A = pi*21^2

Note: the radius in this case is 18+3 = 21 yards (draw out a picture to see this)

A = pi*441

A = 441pi

So the exact area of the Pool+Walkway is 441pi square yards

-------------------------------------------------------

Area of the Pool Only

A = pi*r^2

A = pi*18^2

A = pi*324

A = 324pi

and the area of the Pool only is exactly 324pi square yards.

-------------------------------------------------------

Area of Walkway = (Area of Pool+Walkway) - (Area of Pool Only)

Area of Walkway = (441pi) - (324pi)

Area of Walkway = 441pi - 324pi

Area of Walkway = (441 - 324)pi

Area of Walkway = (117)pi

Area of Walkway = 117pi

So the exact area of the walkway only is 117pi

To find the approximate area, replace pi with 3.14159 (use more digits of pi for greater accuracy) to get

117pi = 117(3.14159) = 367.56603


=====================================================================================
Answers:

Exact area of the walkway only: 117pi square yards

Approximate area of the walkway only: 367.56603 square yards